OpenStudy (anonymous):

Two angles of a triangle have the same measure and the third one is 18 degrees greater than the measure of each of the other two. Find the measure of the LARGEST angle in the triangle. Answer: The LARGEST angle has a measure of ?? degrees.

OpenStudy (ajprincess):

Let x be the measure of one of the angles. Then the other will also measure x. The measure of largest angle= x+18 Sum of the interior angles of a triangle=180 x+x+x+18=180 Can u find the largest angle nw @gil714?

OpenStudy (anonymous):

54

Directrix (directrix):

Is 54 the LARGEST angle?

OpenStudy (ajprincess):

x=54. hw abt the value of the largest angle?

OpenStudy (anonymous):

idk how to get that

OpenStudy (anonymous):

72

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

sorry how about this one .A rectangular room is 3 times as long as it is wide, and its perimeter is 40 meters. Find the dimension of the room. The length is ?? meters and the width is ?? meters.

OpenStudy (ajprincess):

Let the length of the rectangle be l and width of the rectangle be w. l=3w Perimeter=l+l+w+w 40 =2(l+w) 40/2=l+w 20=3w+w Can u find w nw?